OpenStudy (anonymous):

The measure of an angle is twice the measure of its supplement. how many angles are there? What is their relationship? How can you use algebra, such as using the variable x, to help you?

OpenStudy (cwrw238):

supplementary angles add up to 180 degrees if angle is x what is the formula for its supplement?

OpenStudy (anonymous):

x + 2x = 180?

OpenStudy (cwrw238):

supplement of x = 180-x x is twice that so: x = 2(180 - x) solve this to find x

OpenStudy (cwrw238):

can u do that?

OpenStudy (anonymous):

@AccessDenied @princessgoddess10 @scribble @gtaha @kankoo @mumina1999 anyone... help

OpenStudy (anonymous):

yes? i think so

OpenStudy (anonymous):

what

OpenStudy (cwrw238):

ok first distribute 2 over the parentheses

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

gave me a medal

OpenStudy (anonymous):

3x=120?

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

you have facebook

OpenStudy (anonymous):

me?

OpenStudy (cwrw238):

x = 360 - 2x 3x = 360

OpenStudy (anonymous):

gave me a medal

OpenStudy (anonymous):

x=120

OpenStudy (cwrw238):

yea

OpenStudy (anonymous):

What is the height of the triangle if the area is 60 m?